# OpenTCM: A GraphRAG-Empowered LLM-based System for Traditional Chinese Medicine Knowledge Retrieval and Diagnosis

### v3

@@ Abstract
Traditional Chinese Medicine (TCM) represents a rich repository of ancient medical knowledge that continues to play an important role in modern healthcare. Due to the complexity and breadth of the TCM literature, the integration of AI technologies is critical for its modernization and broader accessibility. However, this integration poses considerable challenges, including the interpretation of obscure classical Chinese texts and the modeling of intricate semantic relationships among TCM concepts. In this paper, we develop OpenTCM, an LLM-based system that combines a domain-specific TCM knowledge graph and Graph-based Retrieval-Augmented Generation (GraphRAG). First, we extract more than 3.73 million classical Chinese characters from 68 gynecological books in the Chinese Medical Classics Database, with the help of TCM and gynecology experts. Second, we construct a comprehensive multi-relational knowledge graph comprising more than 48,000 entities and 152,000 interrelationships, using customized prompts and Chinese-oriented LLMs such as DeepSeek and Kimi to ensure high-fidelity semantic understanding. Last, we integrate OpenTCM with this knowledge graph, enabling high-fidelity ingredient knowledge retrieval and diagnostic question-answering without model fine-tuning. Experimental evaluations demonstrate that OpenTCM achieves mean expert scores (MES) of 4.378 in ingredient information retrieval and 4.045 in diagnostic question-answering tasks, outperforming state-of-the-art solutions in real-world TCM use cases.
@@ I Introduction
Traditional Chinese Medicine (TCM) represents a rich repository of ancient medical knowledge that continues to hold significant value in modern healthcare systems [1]. Due to the complexity of the literature and the practice of TCM, many researchers have tried to build systems to assist with consultation and diagnosis for TCM physicians. Current TCM knowledge retrieval and diagnosis systems predominantly often follow a paradigm of fine-tuning general-purpose Large Language Models (LLMs), such as ChatGPT [2] or Claude [3], on domain-specific datasets. However, they often face significant challenges such as the complexity of parsing classical Chinese texts, the intricate relationships among TCM concepts (such as ingredients, symptoms, and diagnostic principles), and the absence of structured knowledge such as knowledge graphs and retrieval-augmented generation (RAG) [4, 5]. In particular, ancient TCM literature is often written in archaic language, making them difficult for modern practitioners and computational models to interpret. Furthermore, existing approaches primarily focus on modern medical literature, overlooking the foundational corpus of classical TCM. Although recent advances in LLMs have proven to be effective in medical knowledge democratization, the direct application of general-purpose LLMs to TCM still faces significant limitations, such as hallucination in prescription generation and semantic understanding of classical Chinese.
@@ III System Overview
In this section, we will describe the system overview of the OpenTCM. As shown in Figure 1, OpenTCM consists of three main components: (1) data collection and pre-processing, (2) knowledge graph construction, and (3) graph-based retrieval-augmented generation.
@@ III-B Knowledge Graph Construction
We use LLMs to extract structured knowledge graphs from the extensive corpus of ancient Chinese characters in TCM literature. However, there are two primary challenges: (1) accurately interpret the semantics of ancient Chinese texts, and then (2) elicit precise, structured responses—such as ingredients, diseases, symptoms, treatments and ingredient references. To overcome these issues, we leverage Chinese-oriented LLMs, which excel in capturing the linguistic and contextual nuances of Chinese literature, and employ carefully crafted prompts to steer the models toward generating detailed and accurate outputs. This methodology ensures the extraction of high-quality information, encompassing symptoms, prescriptions (e.g., ingredient names and dosages), and comprehensive treatment plans. Moreover, the construction of the knowledge graph relies on a systematic prompt-based extraction process tailored to TCM-specific data. Key components of our customized prompt design are as follows.
  • Role Definition: The prompt defines the model as a “TCM Data Processing Assistant,” emphasizing its proficiency in handling classical Chinese texts and its focus on extracting relevant information from TCM literature.
  • Task Description: The prompt instructs the model to extract structured information from the content of TCM chapters and output the results in JSON format. Specifically, the model is required to extract four types of information: (1) book chapter information, including chapter names, IDs, and content, which are stored in a “book” table; (2) treatment plan information, covering treatment names, indications, symptoms, and notes, stored in a “treatment” table; (3) ingredient information, including ingredient names and descriptions, stored in a “ingredient” table; and (4) ingredient-treatment relationship information, which involves ingredient names, dosages, and preparation methods, stored in a “ingredient_ref” table while ensuring the uniqueness of each combination of ingredient ID, treatment ID, and reference ID.
  • Structured Output: The prompt requires the model to output the extracted information in a pre-defined JSON format based on the specified table structures, facilitating subsequent knowledge graph construction and application.
  • Example Input and Output: The prompt includes a sample TCM text and the corresponding expected JSON output to illustrate the task requirements and guide the model to understand the desired extraction and formatting processes.
Through the integration of these models and customized prompts, we have successfully built a knowledge base comprising 3,737 ingredients, 17,031 treatments, and 65,847 ingredient references. The resulting TCM knowledge graph encapsulates 48,406 entities and 152,754 interrelationships, among which 10 types of triplets are formed, as depicted in Table III.
@@
The whole TCM Knowledge Graph and a partial graph are visualized in Figures 2 and 3. This structured knowledge graph not only preserves the richness of TCM data but also facilitates advanced analysis and application in modern research contexts.
@@ III-C Graph-based Retrieval-Augmented Generation
We use GraphRAG [6, 19] to integrate graph-structured data into OpenTCM to enhance retrieval and generation processes. While conventional RAG approaches predominantly processes unstructured text or images, GraphRAG leverages the rich structural relationships embedded within a knowledge graph to improve both information access and reasoning capabilities. This is particularly valuable for TCM, which features complex inter-dependencies among ingredients, symptoms, treatments, and syndromes.
@@ IV-C Real-world Applications
To assess OpenTCM’s effectiveness on real-world applications, we evaluate its performance on two critical downstream tasks: ingredient medicine information retrieval and diagnostic question answering. These tasks test the system’s ability to leverage the TCM knowledge graph and GraphRAG technology to provide accurate, contextually relevant, and user-friendly responses.
@@ IV-C1 Metrics
IV-C2 Ingredient Information Retrieval
@@
Experimental results are presented in Table VI. OpenTCM achieves the highest Mean Expert Score (4.378), surpassing both general-purpose models and TCM-specific baselines. This demonstrates the effectiveness of knowledge graph-enhanced reasoning in improving retrieval precision and contextual relevance.
Figure 5: Score distribution of different models for Ingredient Information Retrieval across four experts. OpenTCM shows significantly higher concentration at Score 4 and 5.
@@ IV-C3 Diagnostic Question Answering
Results in Table VII show that OpenTCM maintains strong performance across all metrics, particularly in interpretability and contextual coherence, outperforming general LLMs and remaining competitive with TCM-specific models.

### v4

@@ Abstract
Traditional Chinese Medicine (TCM) represents a rich repository of ancient medical knowledge that continues to play an important role in modern healthcare. Due to the complexity and breadth of the TCM literature, the integration of AI technologies is critical for its modernization and broader accessibility. However, this integration poses considerable challenges, including the interpretation of obscure classical Chinese texts and the modeling of intricate semantic relationships among TCM concepts. In this paper, we develop OpenTCM, an LLM-based system that combines a domain-specific TCM knowledge graph and Graph-based Retrieval-Augmented Generation (GraphRAG). First, we extract more than 3.73 million classical Chinese characters from 68 gynecological books in the Chinese Medical Classics Database, with the help of TCM and gynecology experts. Second, we construct a comprehensive multi-relational knowledge graph comprising more than 48,000 entities and 152,000 interrelationships, using customized prompts and Chinese-oriented LLMs such as DeepSeek and Kimi to ensure high-fidelity semantic understanding. Last, we empower OpenTCM with GraphRAG, enabling high-fidelity ingredient knowledge retrieval and diagnostic question-answering without model fine-tuning. Experimental evaluations demonstrate that OpenTCM achieves mean expert scores (MES) of 4.378 in ingredient knowledge retrieval and 4.045 in diagnostic question-answering tasks, outperforming state-of-the-art solutions in real-world TCM use cases.
@@ I Introduction
Traditional Chinese Medicine (TCM) represents a rich repository of ancient medical knowledge that continues to hold significant value in modern healthcare systems [1]. Due to the complexity of the literature and the practice of TCM, many researchers have tried to build systems to assist with consultation and diagnosis for TCM physicians [2, 3]. Current TCM knowledge retrieval and diagnosis systems predominantly often follow a paradigm of fine-tuning general-purpose Large Language Models (LLMs), such as Llama [4], on domain-specific datasets. However, they often face significant challenges such as the complexity of parsing classical Chinese texts, the intricate relationships among TCM concepts (such as ingredients, symptoms, and diagnostic principles), and the absence of structured knowledge such as knowledge graphs [5, 6]. In particular, ancient TCM literature is often written in archaic language, making them difficult for modern practitioners and computational models to interpret. Furthermore, existing approaches primarily focus on modern medical literature, overlooking the foundational corpus of classical TCM [7]. Although recent advances in LLMs have proven to be effective in medical knowledge democratization, the direct application of general-purpose LLMs to TCM still faces significant limitations, such as hallucination in prescription generation and semantic understanding of classical Chinese.
@@ III System Overview
In this section, we will describe the system overview of the OpenTCM. As shown in Figure 1, OpenTCM consists of three main components: (1) data collection and pre-processing, (2) knowledge graph construction, and (3) graph-based retrieval-augmented generation. Real-world applications are detailed in section IV.
@@ III-B Knowledge Graph Construction
We use LLMs to extract structured knowledge graphs from a large corpus of classical Chinese medical texts. This process faces two primary challenges: (1) accurately interpreting ancient Chinese medical language and (2) extracting high-quality, structured information such as diseases, symptoms, treatments, and herb references. To address these challenges, we leverage Chinese-centric LLMs and carefully designed prompts. Our final knowledge graph includes 48,406 entities and 152,754 relations, among which 10 types of triplets are formed, as depicted in Table III..
The core of our approach lies in the design of an effective system prompt. Below, we summarize its key components:
  • Role Specification: The prompt defines the model as a “TCM knowledge analysis assistant”, setting expectations for professional, structured, and accurate output.
  • Context Injection: Metadata such as book name, chapter name, and chunk index are included to enhance reference resolution and enable extraction of meta-relations like “Belong to Book”.
  • Task Instructions: The model is explicitly instructed to extract only 10 predefined relation types (e.g., “Treat Disease,” “Use Ingredient”) and return results as a list of (subject, predicate, object) triples.
  • Output Format Constraints: The output must strictly conform to a JSON array of objects, each containing subject, predicate, and object. Invalid or ambiguous content is to be skipped.
  • Few-shot Demonstration: An example passage from a TCM classic and its corresponding JSON-formatted triples are provided to illustrate expected behavior and reinforce output format.
@@
This prompt design effectively guides LLMs to extract structured data from unstructured ancient text, ensuring high consistency, interpretability, and ease of downstream processing for large-scale knowledge graph construction. The whole TCM Knowledge Graph and two partial subgraphs are visualized in Figures 2. This structured knowledge graph not only preserves the richness of TCM data but also facilitates advanced analysis and application in modern research contexts.
@@ III-C Graph-based Retrieval-Augmented Generation
We use GraphRAG [18, 19, 8] to integrate graph-structured data into OpenTCM to enhance retrieval and generation processes. While conventional RAG approaches predominantly processes unstructured text or images, GraphRAG leverages the rich structural relationships embedded within a knowledge graph to improve both information access and reasoning capabilities. This is particularly valuable for TCM, which features complex inter-dependencies among ingredients, symptoms, treatments, and syndromes.
@@ IV-C Real-world Applications
To assess OpenTCM’s effectiveness on real-world applications, we evaluate its performance on two critical downstream tasks: ingredient medicine knowledge retrieval and diagnostic question answering. These tasks test the system’s ability to leverage the TCM knowledge graph and GraphRAG technology to provide accurate, contextually relevant, and user-friendly responses.
@@ IV-C1 Metrics
IV-C2 Ingredient Knowledge Retrieval
@@
Experimental results are presented in Table VI and Figure 4. OpenTCM achieves the highest Mean Expert Score (4.378), surpassing both general-purpose models and TCM-specific baselines. This demonstrates the effectiveness of knowledge graph-enhanced reasoning in improving retrieval precision and contextual relevance.
Figure 4: Score distribution of different models for Ingredient Knowledge Retrieval across four experts. OpenTCM shows significantly higher concentration at Score 4 and 5.
@@ IV-C3 Diagnostic Question Answering
Results in Table VII and Figure 5 show that OpenTCM maintains strong performance across all metrics, particularly in interpretability and contextual coherence, outperforming general LLMs and remaining competitive with TCM-specific models.